QD=50 – 2 P + 0.5 PR and Qs = -4 + P. Here PR if the price of a related good.    If govt. imposes tax on the related good by Re.1 per unit, then how the equilibrium will change?  



1
Expert's answer
2020-09-15T10:10:41-0400

In equilibrium QD = QS, so:

"50 \u2013 2 P + 0.5 PR = -4 + P,"

P = 18 + 1/6PR.

If govt. imposes tax on the related good by Re.1 per unit, then the equilibrium price will increase by 1/6*1 = Re. 0.167.